A house and a piece of land are sold for €850 000.
The house is sold for one and a half times as much
as the land. Find how much the house sold for.
1
Expert's answer
2019-08-13T11:33:45-0400

Let the house be sold for x thousands. Then 2x/3+x=850. Consequently, 5x=3* 850 or x=3*170 = 510, 000.